# DreamStory: Open-Domain Story Visualization by LLM-Guided Multi-Subject Consistent Diffusion

### v2

@@ Abstract
publicationid: pubid: $0$–$00$ 2021 IEEE
@@ 1 Introduction
On the other hand, although previous works [21, 22, 23, 24, 25, 26, 27] have made efforts to improve the consistency in the diffusion model, their attempts have not yielded a satisfactory level of multi-subject consistency in open domain story. These methods fall into four categories: (1) dataset-based training, (2) few-shot fine-tuning, (3) encoder-based, and (4) training-free methods. Dataset-based methods [25] rely on specific story datasets (e.g., PororoSV [28] and FlintstonesSV [29]). Therefore, they are closed-domain methods and are limited in open-domain capabilities. Few-shot fine-tuning methods [23, 24] offer customization but necessitate additional training costs for each story, inevitably leading to overfitting and diversity degradation. Encoder-based methods [22, 30] aim to train an image encoder to convert the reference image into the image condition aligned with the original text condition. It guides the generation process by injecting image conditions into the cross-attention layer. However, these methods mainly focus on a single subject and are hindered by diversity degradation [27] and computational resource consumption. Training-free methods [26, 27, 31] maintain subject consistency by facilitating interaction between target and reference images in the self-attention layer. These methods have gained widespread attention due to their efficiency. However, they still face issues of subject confusion and overlooking fine-grained descriptions in open-domain story visualization.
@@ 2.2.2 LLM in Image Generation
Advanced visual generation models struggle with low-quality descriptions, which impedes their comprehension of subtle semantics. Many research [72, 73, 74, 75, 76, 77] efforts have aimed to enhance the capabilities of T2I models by refining datasets and modifying user prompts. Specifically, Segalis et al. [72] enhanced generation performance by re-captioning the images using a specialized LLM and retraining a text-to-image model on the updated data. Some approaches propose rewriting user prompts to enhance generated images regarding the aesthetic [75] and NFT market values [76]. Yang et al. [78] utilizes language models for planning, recaptioning, and generating images with coherent layouts. Cheng et al. [79] proposes employing LLM to facilitate user editing in an interactive manner.
These methods all indicate that powerful LLM has the potential to enhance image generation. In this article, we utilize the LLM as a director to guide the generation of a series of story images with consistent multi-subject.
@@ 3.1 Overall Framework
    Multimodal anchors for scene generation. The subject portraits, accompanied by their textual descriptions, act as multimodal anchors. The subsequent T2I model leverages these multimodal anchors to maintain subject consistency. It enriches the scenes with additional details, resulting in high-quality visual representations. These images can be transformed into video clips using an Image-to-Video (I2V) model, such as SVD [74], ConsistI2V [80], and Kling  [^1] Kling.
@@ 3.2 LLM Prompt Generation Model
Figure 3: The illustration of our Multi-Subject consistent Diffusion models (MSD), along with its Masked Mutual Self-Attention (MMSA) and Masked Mutual Cross-Attention (MMCA) mechanisms. It uses two subjects as examples and can be extended to any number of subjects. Query, Key, and Value projection in the attention layer have been omitted for ease of presentation.
@@ 3.3.1 Existing Attention Mechanism
A standard attention layer in the popular diffusion model (e.g., SD [6], SD-XL [8], and Playground [9]) can be formulated as follows,
@@
where the $O_{i}$ represents the attention output and $A_{i}$ indicates the attention weight for the i-th image. $Q$ is the query features projected from the spatial features, and $K$ $V$ are the key and value features projected from the spatial features (in self-attention layers) or the textual embedding (in cross-attention layers) with corresponding projection matrices. A simple convolution layer $\textit{conv}_{\scriptsize\tiny out}$ is finally applied to fuse the output features. We omitted the residual connection and layer count to simplify our expression.
Existing work has verified that self-attention can control the appearance of the generated image [85], while cross-attention controls the layout and can be used to locate the area of the target subject [86]. Based on these discoveries, recent works have verified the appearance information of reference images can be injected into the generation process by substiting [26] or cascading [27] $K$ and $V$ with that of in the reference images. In the multi-subject scenario, however, they failed to keep multi-subject consistency because all subjects in the target image can refer to the information from all reference images regardless of whether their roles are the same.
@@ 3.3.2 Accurate Object Mask Generation
Accurate subject mask generation has been verified as a crucial problem in image generation [26, 27] and editing [65, 87]. However, obtaining the subject mask in an unregenerated target image is difficult. Previous works utilize LLM to manage the layout of generated images for editing [79] and accurate attribute binding [78]. This potentially leads to a lack of aesthetic layout and the generation of objects with unreasonable sizes, as LLM has not been optimized in this situation. Since diffusion models tend to generate similar layouts under close control conditions if the random seeds are the same [88], we adopt an open-vocabulary segmentation model, e.g., GroundingSAM [89], to obtain an accurate subject mask in rehearsal target images, which is pre-generate with the original diffusion model. The detection phrases are marked by LLM, as mentioned above. To improve accuracy, we contact tokens of all subjects separated by periods as detection prompts for the target image, such as "man. girl.". A simple post-processing is adopted to guarantee the non-overlapping of all masks, enhancing the robustness of our approach.
@@
Due to the imposition of a new control process, the image may go beyond the SAM’s mask during generation, especially in the later steps. So, it is necessary to adjust the mask according to the features in the generation process. Previous work [26, 27] mainly adopts a segmentation mask by averaging the cross-attention maps of the subject token. However, this strategy may create a holed and noisy mask [87]. Therefore, we obtain a segmentation mask by multiplying the self-attention and cross-attention maps, with the self-attention map serving as a completion of the cross-attention maps [87]. So, the semantics maps are calculated as follows,
$m_{i}$ $=\textit{mean}(\sum_{l\in L}\sum_{r\in R}(A_{sa})^{r}\times A_{ca}),$ (3)
$M_{i}$ $=\textit{Otsu}\Big{(}F(m_{\text{\scriptsize\tiny TGT}}^{i})\Big{)}^{\top}\times\textit{Otsu}\Big{(}F(m_{i})\Big{)},i\neq\text{\scriptsize TGT}$ (4)
@@
where $A_{sa}$ and $A_{ca}$ denote the self-attention and cross-attention layers in the same block. The mask of the target image, denoted with the subscript ‘TGT’. $L$ denotes the layer for gathering the attention map, and the averaging operation is denoted by $\textit{mean}(\cdot)$. $R$ is a hyper-parameter set as 4 followed by [87]. To save computational cost, we only collect all layers $L$ from the previous timestep $t-1$ to calculate the mask. The threshold is determined by Otsu’s method [90], represented as $\textit{Otsu}(\cdot)$. The flatten operation, $F(\cdot)$, converts the matrix into a one-dimensional array. The symbol $\times$ stands for matrix multiplication. The final matrix $M_{i}$ illustrates the correlation between the elements of the target and reference images, ensuring that elements in the target image reference only the related regions in the reference image. It should be noted that this is an optional strategy, employed only when the re-generated targets significantly exceed the initial range.
@@ 3.3.3 Masked Mutual Self-Attention
where $M_{i}$ is the subject mask for i-th reference images, and $\oplus$ indicates the concatenation operation. We assume the last one to be the target image, denoted with the subscript ‘TGT’. The standard attention masking technique is adopted, which nullifies softmax’s logits by assigning their scores to $\infty$- ∞ based on the mask, followed by previous works [26, 27]. It should be noted that, unlike ConsiStory [27] and StoryDiffusion [31], which allows all areas of the target image to reference the subject in the reference image, our method only permits referencing information from the same subject.
@@ 3.3.4 Masked Mutual Cross-Attention
$A_{\text{\scriptsize\tiny TGT}}^{i}$ $=\textit{softmax}\left({Q_{\text{\scriptsize\tiny TGT}}K_{i}}/{\sqrt{d_{k}}}+\log(F(m_{\text{\scriptsize\tiny TGT}}^{i})\times 1)\right),$ (10)
@@
Figure 4: Qualitative comparisons of our DreamStory with SOTA approaches on the FSS real story benchmark. Ours, MuDI, and ConsiStory utilize the subject image on the bottom-left as the reference image. In contrast, StoryDiffusion references the subject image on the bottom-right. Different subjects are indicated with different colors. Please refer to the supplementary materials to watch the video.
Figure 5: Qualitative comparisons of our DreamStory with SOTA approaches on the ChatGPT generated story benchmark. Ours, MuDI, and ConsiStory utilize the subject image on the bottom-left as the reference image. In contrast, StoryDiffusion references the subject image on the bottom-right. Different subjects are indicated with different colors. Please refer to the supplementary materials to watch the video.
@@ 4 Experiments
This section will introduce the evaluation benchmark and metrics, implementation details, and comprehensive experimental results. In Sec. 4.1, we will introduce the constructed benchmark, which includes 100 stories and 400 synthetic cases. Sec. 4.2 will introduce the objective and subjective evaluation metrics. We present the specific implementation details in Sec. 4.3, including our LLM, diffusion backbone, and MSD module. In Sec. 4.4 and Sec. 4.5, we will respectively present the comparative results with the current state-of-the-art (SOTA) methods and conduct an ablation study. Sec. 4.7 briefly discuss the limitations of our method.
@@ 4.1.2 The 400 Synthetic Benchmark
TABLE I: Quantitative results of different backbone for our DreamStory on the DS-500 benchmark. Red indicate the best performance.
@@ 4.2.2 Subjective Evaluation
Due to the bias of existing metrics, a user study is conducted to assess subjective results. Given the variability of individual ratings and the broad spectrum of scores across different evaluators, we employed a pairwise comparison in our user study. For each evaluation, two sets of images were randomly displayed, each generated by a different method and accompanied by their respective texts. Participants were asked to judge each metric by selecting one of three options: Image A is superior, Image B is superior, or both are comparable. We engaged 20 independent evaluators for the assessment. Each evaluator conducted 100 reviews per benchmark, culminating in 2000 votes in total. The final results were compiled and are presented as percentages.
Figure 7: Ablation studies of different generation results. All settings except the baseline utilize the subject image on the left as the reference image. Different subjects are indicated with different colors. Our method better maintains consistency across multiple subjects, such as the woman and cat in the first row, the pirate and Siamese cat in the second row, the pirate and cat in the third row, and the boy in the last row.
@@
Figure 8: Qualitative comparisons of our DreamStory with StoryGen on their benchmark. All approaches utilize the subject image on the left as the reference image. Different subjects are indicated with different colors. The narrative text of the story is presented below and serves as the input for our DreamStory. Please refer to the supplementary materials to watch the video.
Figure 9: Qualitative comparisons of our DreamStory with TheaterGen on their benchmark. Different subjects are indicated with different colors. The narrative text of the story is presented on the left and serves as the input for our DreamStory. Please refer to the supplementary materials to watch the video.
Figure 10: Failure Cases of LLM.
@@ 4.3 Implementation Details
Attention Mechanisms in MSD. Our MSD is applied across all diffusion steps to ensure multi-subject consistency. The masked mutual self-attention is applied to all decoder layers to maintain the appearance consistency, followed by previous works [26, 27]. Inspired by previous work [30, 88], the masked mutual cross-attention is applied to all layers for better cross-attention fusion. The dropout [27] strategy with a dropout rate of 0.5 is adopted to enhance layout diversity. Furthermore, we adopt the open-vocabulary segmentation model, GroundingSAM [89], to generate precise masks for the subjects. This process begins with the detection of the subject using the open-vocabulary detection model, GroundingDINO [92], followed by segmentation with the powerful SAM [95].
@@ 4.4.1 Objective Comparison
The overall results are presented in Tab. II. As can be seen from the table, our DreamStory outperforms other methods in terms of all metrics. Notably, the D&C-DS metric of ours is significantly surpassed other methods, exceeding MuDI, ConsiStory, and StoryDiffusion by margins of 0.1034(23.4%), 0.1293(25.1%), and 0.1080(24.7%) respectively in the 2-Subject of DS-500 benchmark. This pattern is mirrored in the 3-Subject benchmark, reinforcing the effectiveness of our method in maintaining multi-subject consistency. Furthermore, our method exhibits a notable advantage on the DS metric, outperforming other training-free SOTA methods (ConsiStory and StoryDiffusion) on the 2-Subject benchmark by at least 0.05 (9.0%). This lead, however, narrows to an approximate average of 0.02 (3.9%) on the 3-Subject benchmark. This is attributed to the limitations of the diffusion model when generating three subjects, leading to a higher likelihood of subject fusion, which is also discussed in MuDI [24].

### v3

@@ Abstract
publicationid: pubid: 0000–0000/00\$00.00 © 2021 IEEE
@@ 1 Introduction
On the other hand, although previous works [21, 22, 23, 24, 25, 26, 27] have made efforts to improve the consistency in the diffusion model, their attempts have not yielded a satisfactory level of multi-subject consistency in open-domain stories. These methods fall into four categories: (1) dataset-based training, (2) few-shot fine-tuning, (3) encoder-based, and (4) training-free methods. Dataset-based methods [25] rely on specific story datasets (e.g., PororoSV [28] and FlintstonesSV [29]). Therefore, they are closed-domain methods and are limited in open-domain capabilities. Few-shot fine-tuning methods [23, 24] offer customization but necessitate additional training costs for each story, inevitably leading to overfitting and diversity degradation. Encoder-based methods [22, 30] aims to train an image encoder to convert the reference image into the image condition aligned with the original text condition. It guides the generation process by injecting image conditions into the cross-attention layer. However, these methods mainly focus on a single subject and are hindered by diversity degradation [27] and computational resource consumption. Training-free methods [26, 27, 31] maintain subject consistency by facilitating interaction between target and reference images in the self-attention layer. These methods have gained widespread attention due to their efficiency. However, they still face issues of subject confusion and overlooking fine-grained descriptions in open-domain story visualization.
@@ 2.2.2 LLM in Image Generation
Advanced visual generation models struggle with low-quality descriptions, which impedes their comprehension of subtle semantics. Many research [72, 73, 74, 75, 76, 77] efforts have aimed to enhance the capabilities of T2I models by refining datasets and modifying user prompts. Specifically, Segalis et al. [72] enhanced generation performance by re-captioning the images using a specialized LLM and retraining a text-to-image model on the updated data. Some approaches propose rewriting user prompts to enhance generated images regarding the aesthetic [75] and NFT market values [76]. Yang et al. [78] utilizes language models for planning, recaptioning, and generating images with coherent layouts. Cheng et al. [79] proposes employing LLM to facilitate user editing in an interactive manner. Zhuang et al. [80] leverages an LLM-guided pipeline to generate minute-long vlogs with coherent storylines and diverse scenes.
These methods all indicate that a powerful LLM has the potential to enhance image generation. In this article, we utilize the LLM as a director to guide the generation of a series of story images with consistent multi-subject.
@@ 3.1 Overall Framework
    Multimodal anchors for scene generation. The subject portraits, accompanied by their textual descriptions, act as multimodal anchors. The subsequent T2I model leverages these multimodal anchors to maintain subject consistency. It enriches the scenes with additional details, resulting in high-quality visual representations. These images can be transformed into video clips using an Image-to-Video (I2V) model, such as SVD [74], ConsistI2V [81], and Kling  [^1] Kling, or extended into longer videos with models like SEINE [82].
@@ 3.2 LLM Prompt Generation Model
Figure 3: The illustration of our Multi-Subject consistent Diffusion models (MSD), along with its Masked Mutual Self-Attention (MMSA) and Masked Mutual Cross-Attention (MMCA) mechanisms. It uses two subjects as examples and can be extended to any number of subjects. Query, Key, and Value projections in the attention layer have been omitted for ease of presentation.
@@ 3.3.1 Existing Attention Mechanism
Assume the resolution of the generated image is downsampled to $H\times W$ in the attention layer, and $B$ and $C$ respectively denote the batch size and the number of channels. A standard attention layer in the popular diffusion model (e.g., SD [6], SD-XL [8], and Playground [9]) can be formulated as follows,
@@
where the $O_{i}\in\mathbb{R}^{\footnotesize{B\times C\times HW}}$ represents the attention output and $A_{i}\in\mathbb{R}^{\footnotesize{BC\times HW\times S}}$ indicates the attention weight for the i-th image, with $S$ being the embedding length of key-value features. For self-attention, the key-value features $KV$ come from spatial features with embedding length $S=HW$, while for cross-attention, they are derived from text embeddings where $S=T=77$ in U-Net-based models [6, 8, 9]. The query features $Q$ are always projected from spatial features through learned projection matrices. A simple convolution layer $\textit{conv}_{\scriptsize\tiny out}$ is finally applied to fuse the output features. We omitted the residual connection and layer count to simplify our expression.
Existing work has verified that self-attention can control the appearance of the generated image [87], while cross-attention controls the layout and can be used to locate the area of the target subject [88]. Based on these discoveries, recent works have verified that the appearance information of reference images can be injected into the generation process by substiting [26] or cascading [27] $K$ and $V$ with that of in the reference images. In the multi-subject scenario, however, they failed to keep multi-subject consistency because all subjects in the target image can refer to the information from all reference images regardless of whether their roles are the same.
@@ 3.3.2 Accurate Object Mask Generation
Accurate subject mask generation has been verified as a crucial problem in image generation [26, 27] and editing [65, 89]. However, obtaining the subject mask in an unregenerated target image is difficult. Previous works utilize LLM to manage the layout of generated images for editing [79] and accurate attribute binding [78]. This potentially leads to a lack of aesthetic layout and the generation of objects with unreasonable sizes, as LLM has not been optimized in this situation. Since diffusion models tend to generate similar layouts under close control conditions if the random seeds are the same [90], we adopt an open-vocabulary segmentation model, e.g., GroundingSAM [91], to obtain an accurate subject mask in rehearsal target images, which is pre-generated with the original diffusion model. The detection phrases are marked by LLM, as mentioned above. To improve accuracy, we contact tokens of all subjects separated by periods as detection prompts for the target image, such as "man. girl.". A simple post-processing is adopted to guarantee the non-overlapping of all masks, enhancing the robustness of our approach.
@@
Due to the imposition of a new control process, the image may go beyond the SAM’s mask during generation, especially in the later steps. So, it is necessary to adjust the mask according to the features in the generation process. Previous work [26, 27] mainly adopts a segmentation mask by averaging the cross-attention maps of the subject token. However, this strategy may create a hole and a noisy mask [89]. Therefore, we obtain a segmentation mask by multiplying the self-attention and cross-attention maps, with the self-attention map serving as a completion of the cross-attention maps [89]. So, the semantics maps are calculated as follows,
$m_{\text{\scriptsize\tiny TGT}}^{i}$ $=\textit{mean}(\sum_{l\in L}\sum_{r\in R}(A_{sa})^{r}\times A_{ca}),$ (3)
$M_{i}$ $=\textit{Otsu}\Big{(}m_{\text{\scriptsize\tiny TGT}}^{i}\Big{)}\times\textit{Otsu}\Big{(}m_{\text{\scriptsize\tiny REF}}^{i}\Big{)}^{\top},$ (4)
@@
where $A_{sa}\in\mathbb{R}^{\footnotesize{BC\times HW\times HW}}$ and $A_{ca}\in\mathbb{R}^{\footnotesize{BC\times HW\times T}}$ denote the self-attention and cross-attention layers of the target image in the same block. The flattened mask of the reference image and target image $m^{i}\in\mathbb{R}^{\footnotesize{B\times HW}}$ are denoted with the subscript ‘REF’ and ‘TGT’, respectively. $L$ denotes the layer for gathering the attention map, and the averaging operation is denoted by $\textit{mean}(\cdot)$. $R$ is a hyper-parameter set as 4 followed by [89]. To save computational cost, we only collect all layers $L$ from the previous timestep t1t-1 to calculate the mask. The threshold is determined by Otsu’s method [92], represented as $\textit{Otsu}(\cdot)$. The symbol $\times$ stands for matrix multiplication. The final matrix $M_{i}\in\mathbb{R}^{\footnotesize{B\times HW\times HW}}$ illustrates the correlation between the elements of the target and reference images, ensuring that elements in the target image reference only the related regions in the reference image. It should be noted that this is an optional strategy, employed only when the re-generated targets significantly exceed the initial range.
@@ 3.3.3 Masked Mutual Self-Attention
where $M_{i}$ is the subject mask for i-th reference images, and $\oplus$ indicates the concatenation operation. We assume the last one to be the target image, denoted with the subscript ‘TGT’. The standard attention masking technique is adopted, which nullifies softmax’s logits by assigning their scores to -$\infty$ based on the mask, followed by previous works [26, 27]. It should be noted that, unlike ConsiStory [27] and StoryDiffusion [31], which allow all areas of the target image to reference the subject in the reference image, our method only permits referencing information from the same subject.
@@ 3.3.4 Masked Mutual Cross-Attention
$A_{\text{\scriptsize\tiny TGT}}^{i}$ $=\textit{softmax}\left({Q_{\text{\scriptsize\tiny TGT}}K_{i}}/{\sqrt{d_{k}}}+\log(m_{\text{\scriptsize\tiny TGT}}^{i}\times\mathds{1})\right),$ (10)
@@
Figure 4: Qualitative comparisons of our DreamStory with SOTA approaches on the FSS real story benchmark. Ours, MuDI, and ConsiStory utilize the subject image on the bottom-left as the reference image. In contrast, StoryDiffusion references the subject image on the bottom-right. Different subjects are indicated with different colors. Please visit the project homepage to watch the video.
Figure 5: Qualitative comparisons of our DreamStory with SOTA approaches on the ChatGPT generated story benchmark. Ours, MuDI, and ConsiStory utilize the subject image on the bottom-left as the reference image. In contrast, StoryDiffusion references the subject image on the bottom-right. Different subjects are indicated with different colors. Please visit the project homepage to watch the video.
@@ 4 Experiments
This section will introduce the evaluation benchmark and metrics, implementation details, and comprehensive experimental results. In Sec. 4.1, we will introduce the constructed benchmark, which includes 100 stories and 400 synthetic cases. Sec. 4.2 will introduce the objective and subjective evaluation metrics. We present the specific implementation details in Sec. 4.3, including our LLM, diffusion backbone, and MSD module. In Sec. 4.4 and Sec. 4.5, we will respectively present the comparative results with the current state-of-the-art (SOTA) methods and conduct an ablation study. Sec. 4.7 briefly discusses the limitations of our method.
@@ 4.1.2 The 400 Synthetic Benchmark
TABLE I: Quantitative results of different backbones for our DreamStory on the DS-500 benchmark. Red indicate the best performance.
@@ 4.2.2 Subjective Evaluation
Due to the bias of existing metrics, a user study is conducted to assess subjective results. Given the variability of individual ratings and the broad spectrum of scores across different evaluators, we employed a pairwise comparison in our user study. For each evaluation, two sets of images were randomly displayed, each generated by a different method and accompanied by its respective text. Participants were asked to judge each metric by selecting one of three options: Image A is superior, Image B is superior, or both are comparable. We engaged 20 independent evaluators for the assessment. Each evaluator conducted 100 reviews per benchmark, culminating in 2000 votes in total. The final results were compiled and are presented as percentages.
Figure 7: Ablation studies of different generations results. All settings except the baseline utilize the subject image on the left as the reference image. Different subjects are indicated with different colors. Our method better maintains consistency across multiple subjects, such as the woman and cat in the first row, the pirate and Siamese cat in the second row, the pirate and cat in the third row, and the boy in the last row.
@@
Figure 8: Qualitative comparisons of our DreamStory with StoryGen on their benchmark. All approaches utilize the subject image on the left as the reference image. Different subjects are indicated with different colors. The narrative text of the story is presented below and serves as the input for our DreamStory. Please visit the project homepage to watch the video.
Figure 9: Qualitative comparisons of our DreamStory with TheaterGen on their benchmark. Different subjects are indicated with different colors. The narrative text of the story is presented on the left and serves as the input for our DreamStory. Please visit the project homepage to watch the video.
Figure 10: Qualitative comparison of applying MMSA to U-Net encoder versus decoder layers. Different subjects are indicated with different colors. The results demonstrate that only applying MMSA to decoder layers preserves appearance consistency.
@@ 4.3 Implementation Details
Attention Mechanisms in MSD. Our MSD is applied across all diffusion steps to ensure multi-subject consistency. The Masked Mutual Self-Attention (MMSA) is applied to all decoder layers to maintain the appearance consistency. As shown in Fig. 10, this strategy more effectively preserves the appearance consistency compared to applying MMSA to encoder layers. These observations align with prior works [26, 27], which highlight that the U-Net’s decoder is specialized to handle fine-grained visual details (e.g., colors and textures), while the encoder manages structural information. Inspired by previous work [30, 90], the Masked Mutual Cross-Attention (MMCA) is applied to all layers for better cross-attention fusion. The dropout [27] strategy with a dropout rate of 0.5 is adopted to enhance layout diversity. Furthermore, we adopt the open-vocabulary segmentation model, GroundingSAM [91], to generate precise masks for the subjects. This process begins with the detection of the subject using the open-vocabulary detection model, GroundingDINO [94], followed by segmentation with the powerful SAM [97].
@@ 4.4.1 Objective Comparison
The overall results are presented in Tab. II. As can be seen from the table, our DreamStory outperforms other methods in terms of all metrics. Notably, the D&C-DS metric of ours significantly surpassed other methods, exceeding MuDI, ConsiStory, and StoryDiffusion by margins of 0.1034(23.4%), 0.1293(25.1%), and 0.1080(24.7%) respectively in the 2-Subject of DS-500 benchmark. This pattern is mirrored in the 3-Subject benchmark, reinforcing the effectiveness of our method in maintaining multi-subject consistency. Furthermore, our method exhibits a notable advantage on the DS metric, outperforming other training-free SOTA methods (ConsiStory and StoryDiffusion) on the 2-Subject benchmark by at least 0.05 (9.0%). This lead, however, narrows to an approximate average of 0.02 (3.9%) on the 3-Subject benchmark. This is attributed to the limitations of the diffusion model when generating three subjects, leading to a higher likelihood of subject fusion, which is also discussed in MuDI [24].
@@
Figure 15: Ablation study of accurate mask generation. Different subjects are indicated with different colors. Without our accurate mask generation strategy, the initial mask (left) fails to fully cover the boy, causing inconsistent hair color (highlighted in red circle).